# Highest weight category structures on Rep(B) and full exceptional collections on generalized flag varieties over $\mathbb{Z}$

### v2

@@ 2.2 Joseph-Demazure modules
Let $H\subset G$ be a flat subgroup scheme of $G$. Associated to it are the restriction and induction functors ${{\mathrm{res}}}_{H}^{G}$ and ${{\mathrm{ind}}}_{H}^{G}$, [29, I, Section 3.1]. Recall the definition of Joseph’s functors ${\sf H}_{w},w\in W$, following [46].
Definition 1.
Let $M$ be a $B$-module and $w\in W$. Consider the Schubert variety $X_{w}$ associated to $w$. Then the functor ${\sf H}_{w}:{{\mathrm{Rep}}}(B)\rightarrow{{\mathrm{Rep}}}(B)$ is given by $M\rightarrow{\rm H}^{0}(X_{w},\mathcal{L}(M))$.
Equivalently, the functor ${\sf H}_{w}$ can be defined as in [47, Definition 7.2.1]. Consider the minimal parabolic subgroup ${P}_{\alpha}$ generated by $B$ and ${U}_{\alpha}$. Then
${\sf H}_{w}:={{\mathrm{res}}}_{B}^{{P}_{\alpha_{1}}}{{\mathrm{ind}}}_{B}^{{P}_{\alpha_{1}}}\dots{\mathrm{res}}_{B}^{{P}_{\alpha_{n}}}{{\mathrm{ind}}}_{B}^{{P}_{\alpha_{n}}},$
where $s_{\alpha_{1}}s_{\alpha_{2}}\dots s_{\alpha_{n}}$ is a reduced expression for $w$.
The functor ${\sf H}_{w}$ is left exact; let ${\rm R}{\sf H}_{w}$ denote the right derived functor $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{+}({{\mathrm{Rep}}}(B))\rightarrow\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{+}({{\mathrm{Rep}}}(B))$. It restricts to a functor between bounded derived categories $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({{\mathrm{rep}}}(B))\rightarrow\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({{\mathrm{rep}}}(B))$.
Definition 2 (Dual Joseph modules $P(\lambda)$).
Let $\lambda\in X(T)$. Let $\lambda^{+}$ be the dominant weight in the Weyl group orbit of $\lambda$ and let $w$ be minimal so that $\lambda=w\lambda^{+}$. The dual Joseph module $P(\lambda)$ is set to be ${\sf H}_{w}(\lambda^{+})$.
@@ 2.3 Relative Schubert modules
Definition 3 (Relative Schubert modules $Q(\lambda)$).
Let $\lambda=w\lambda^{+}$ as above. The relative Schubert module $Q(\lambda)$ is set to be the kernel of the (surjective) restriction homomorphism $P(\lambda)\rightarrow{\rm H}^{0}(\partial X_{w},\mathcal{L}(\lambda^{+}))$.
2.4 $B$-cohomological duality
@@
${\mathrm{Ext}}^{i}_{B}(P(\lambda)^{\ast},Q(\mu))=\left\{\begin{array}[]{l}{\sf k},\quad{\rm for}\quad i=0,\ \lambda=-\mu,\\ 0,\quad{\rm otherwise}.\\ \end{array}\right.$ (3)
See also 2.4.
@@ 2.5 Excellent order
Recall that we have fixed a Weyl group invariant inner product (,)(~{},~{})( , ) on $X({T})\otimes_{\mathbb{Z}}\mathbb{R}$.
Definition 4.
([46]) Let $\lambda,\mu\in X({T})$. Define that $\lambda$ is less than $\mu$ in the excellent order, notation $\lambda\leq_{e}\mu$, if either $(\lambda,\lambda)<(\mu,\mu)$ or $\lambda=w\nu$, $\mu=z\nu$ for some $\nu\in X({T})_{+}$, $w,z\in W$ with $w\leq z$ (in the Bruhat order on $W$).
@@ 2.6 Antipodal excellent order
2.7 Highest weight category structures on ${\mathrm{Rep}}(B)$
@@
([46, Theorem 1.6, (i)]) The category ${\mathrm{Rep}}(B)$ of rational representations of $B$ is a highest weight category with respect to the excellent order. The $P(\lambda)$ are the costandard modules for this order.
@@
([46, Theorem 1.6, (ii)]) The category ${\mathrm{Rep}}(B)$ of rational representations of $B$ is a highest weight category with respect to the antipodal excellent order. The $Q(\lambda)$ are the costandard modules for this order.
@@
Let ${\mathrm{Rep}}(B)_{\leq_{a}\lambda}$ denote the subcategory of $B$-modules whose weights $\mu$ all satisfy $\mu\leq_{a}\lambda$. Then, in the usual terminology, cf. [19, §2], the costandard module $Q(\lambda)$ for $\leq_{a}$ is the injective hull of ${\sf k}_{\lambda}$ in ${\mathrm{Rep}}(B)_{\leq_{a}\lambda}$ and the standard module $P(-\lambda)^{*}$ for $\leq_{a}$ is the projective cover of ${\sf k}_{\lambda}$ in ${\mathrm{Rep}}(B)_{\leq_{a}\lambda}$.
2.8 Cohomology vanishing for $B$-modules
@@
[47, Theorem 3.2.6] Let $\lambda,\mu\in X({T})$. Then $H^{p}(B,P(\lambda)\otimes Q(\mu))=0$ for $p>0$.
@@
Let $\lambda,\mu\in X(T)$, $\nu\in X(T)_{+}$. Then $H^{p}(B,P(\lambda)\otimes Q(\mu)\otimes\nabla_{\nu})=0$ for $p>0$.
@@
The vanishing statement of Corollary 2.5 exhibits a remarkable interplay among the three highest weight category structures: one on the category ${\mathrm{Rep}}(G)$ from Section 2.1, and the other two on ${\mathrm{Rep}}(B)$ from Section 2.7. Note also that ${\sf k}=P(0)=Q(0)=\nabla(0)$ is costandard in all three highest weight category structures.
2.9 Filtrations on $B$-modules
We recall here the fundamental statements concerning filtrations on $B$-modules. As we emphasized in the Notation section, one has to take some care when citing the literature, because our $B$ is not $B^{+}$. We have to convert to our conventions. That means that dominant often becomes antidominant and vice versa.
We begin with recalling the definition of good filtration on a $G$-module, [17].
Definition 6.
A rational $G$-module $M$ is said to admit a good filtration provided that there exists a finite or infinite increasing filtration
such that each $M_{i}/M_{i-1}=\nabla_{\lambda_{i}}$ for some $\lambda_{i}\in X(T)_{+}$.
@@ 2.9 Filtrations on $B$-modules
An unfortunate side effect of this definition is that if $M$ admits a good filtration, its dimension is at most countable, even if $G$ acts trivially on $M$. This makes results like Theorem 2.16 a bit complicated. If one wants to keep representation theory clean, it is better to use [47, Definition 4.1.1], which is similar to the next definition.
Definition 7.
[47, Definition 2.3.6] A $B$-module $M$ is said to have an excellent filtration if and only if there exists a filtration $0\subset F_{0}\subset F_{1}\subset\dots$ by $B$-modules such that $\cup F_{i}=M$ and $F_{i}/F_{i-1}=\oplus P(\lambda_{i})$ for some $\lambda_{i}\in X(T)$. Here $\oplus$ stands for any number of copies, ranging from zero copies to infinitely many.
Remark 2.8.
Definition 8.
[47, Definition 2.3.8] A $B$-module $M$ is said to have a relative Schubert filtration if and only if there exists a filtration $0\subset F_{0}\subset F_{1}\subset\dots$ by $B$-modules such that $\cup F_{i}=M$ and $F_{i}/F_{i-1}=\oplus Q(\lambda_{i})$ for some $\lambda_{i}\in X(T)$. Here $\oplus$ stands for any number of copies, ranging from zero copies to infinitely many.
Remark 2.9.
Definition 9.
A weight $\lambda$ of a $B$-module $M$ is called extremal if all other weights $\mu$ of $M$ are shorter or equally long: $(\mu,\mu)\leq(\lambda,\lambda)$.
Proposition 2.10.
[47, Proposition 2.3.11] For any $B$-invariant closed subset $S\subset G/B$ and $\lambda\in X(T)_{+}$, the module $H^{0}(S,\mathcal{L}(\lambda))$ has a relative Schubert filtration with layers $Q(w\lambda)$. Moreover, $Q(w\lambda)$ occurs if and only if $w\lambda$ is an extremal weight of $H^{0}(S,\mathcal{L}(\lambda))$ and then has multiplicity one.
Corollary 2.11.
[47, Corollary 2.3.12] The modules ${\sf H}_{w}(\lambda)$ have relative Schubert filtration for all $w\in W$ and $\lambda\in X(T)$.
Theorem 2.12.
[47, Corollary 5.2.3], [36, §5, Corollary 2] Let $\lambda\in X(T)_{+}$ be a regular dominant weight and $\mu\in X(T)$ be an arbitrary weight. Then $\lambda\otimes Q(\mu)$ has excellent filtration.
Theorem 2.13.
[47, Corollary 5.2.7] Let $\lambda\in X({T})$ and $\mu\in X({T})_{+}$, so that $P(w_{0}\mu)=\nabla_{\mu}$. Then $P(\lambda)\otimes\nabla_{\mu}$ has excellent filtration.
Theorem 2.14.
[35, Theorem 2] Let ${\sf k}$ be a field of characteristic zero. Let $n$ be an integer and $M_{1},\dots M_{n}$ be $n$ weak modules (i.e. each $M_{i}$ has a filtration by $Q(\lambda_{i})$’s, $i=1,\dots,n)$. Suppose $\lambda-n\rho$ is dominant. Then the $B$-module $M_{1}\otimes\dots\otimes M_{n}\otimes{\sf k}_{\lambda}$ is strong (i.e. has a filtration by dual Joseph modules $P(\lambda)$).
2.10 Cohomology criteria for $G$-modules
Theorem 2.15.
[17, Theorem 1.1’] If $\lambda\in X(T)_{+}$ and $M$ is a $G$-module with a good filtration then ${\rm H}^{i}(G,\nabla_{\lambda}\otimes M)=0$ for $i>0$.
Theorem 2.16.
[24, Theorem 4, Corollary 7] Let $M$ be a finitely/countably generated rational $G$-module satisfying ${\rm H}^{1}(G,\nabla_{\mu}\otimes M)=0$ for all $\mu\in X(T)_{+}$. Then $M$ admits a good filtration.
Theorem 2.17.
[29, II Proposition 4.21, B.9] Let $\lambda,\mu\in X(T)_{+}$. Then $\nabla_{\lambda}\otimes\nabla_{\mu}$ admits a good filtration.
2.11 Cohomology criteria for $B$-modules
Theorem 2.18.
[46, Theorem 3.2.7] Let $M$ be a $B$-module such that for every $\lambda\in X(T)$, $H^{1}(B,M\otimes Q(\lambda))=0$. Then, $M$ has an excellent filtration.
Remark 2.19.
For the results over ${\mathbb{Z}}$ we will use ${\mathbb{Z}}$-forms $P(\lambda)_{\mathbb{Z}}$, $Q(\mu)_{\mathbb{Z}}$ of $P(\lambda)$, $Q(\mu)$ respectively. They are constructed and studied in [47, Chapter 7]. Both $P(\lambda)_{\mathbb{Z}}$ and $Q(\mu)_{\mathbb{Z}}$ are finitely generated and free over ${\mathbb{Z}}$. They share many properties with their counterparts over fields. We often drop the ${\mathbb{Z}}$ from the notations $P(\lambda)_{\mathbb{Z}}$, $Q(\mu)_{\mathbb{Z}}$.
3 $T$-equivariant ${\sf K}$-theory of $G/B$

### v3

@@ 2.2 Joseph-Demazure modules
Let $\mathbf{H}\subset\mathbf{G}$ be a flat subgroup scheme of $\mathbf{G}$. Associated to it are the restriction and induction functors ${{\mathrm{res}}}_{\mathbf{H}}^{\mathbf{G}}$ and ${{\mathrm{ind}}}_{\mathbf{H}}^{\mathbf{G}}$, [32, I, Section 3.1]. Recall the definition of Joseph’s functors ${{\sf H}}_{w},w\in W$, following [48].
Definition 2.1.
Let $M$ be a $\mathbf{B}$-module and $w\in W$. Consider the Schubert variety ${X}_{w}=\overline{\mathbf{B}w\mathbf{B}}/\mathbf{B}$ associated to $w$. Then the functor ${{\sf H}}_{w}:{{\mathrm{Rep}}}(\mathbf{B})\rightarrow{{\mathrm{Rep}}}(\mathbf{B})$ is given by $M\rightarrow{\rm H}^{0}({X}_{w},\mathcal{L}(M))$.
Equivalently, by [49, Proposition 2.2.5], the functor ${{\sf H}}_{w}$ can be described as follows. If $\alpha\in\Pi$ with corresponding simple reflection $s$, let $\mathbf{P}_{s}$ or $\mathbf{P}_{\alpha}$ denote the minimal parabolic subgroup generated by $\mathbf{B}$ and ${\mathbf{U}}_{\alpha}$. Then ${\sf H}_{s}={{\mathrm{res}}}_{\mathbf{B}}^{{\mathbf{P}}_{s}}{{\mathrm{ind}}}_{\mathbf{B}}^{{\mathbf{P}}_{s}}$. And if $s_{1}s_{2}\dots s_{n}$ is a reduced expression for $w$, then
${{\sf H}}_{w}:={\sf H}_{s_{1}}\circ\dots\circ{\sf H}_{s_{n}}.$
The functor ${{\sf H}}_{w}$ is left exact; let ${\rm R}{{\sf H}}_{w}$ denote the right derived functor $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{+}({{\mathrm{Rep}}}(\mathbf{B}))\rightarrow\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{+}({{\mathrm{Rep}}}(\mathbf{B}))$. It restricts to a functor between bounded derived categories $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({{\mathrm{rep}}}(\mathbf{B}))\rightarrow\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({{\mathrm{rep}}}(\mathbf{B}))$ (see Section 8.2).
Definition 2.2.
(Dual Joseph modules $P(\lambda)$). Let $\lambda\in X(\mathbf{T})$. Let $\lambda^{+}$ be the dominant weight in the Weyl group orbit of $\lambda$ and let $w$ be minimal so that $\lambda=w\lambda^{+}$. The dual Joseph module $P(\lambda)$ is set to be ${{\sf H}}_{w}(\lambda^{+})$. Its $\mathbf{B}$-socle is of weight $\lambda$ [49, Lemma 2.2.9].
@@ 2.3 Relative Schubert modules
Definition 2.3 (Relative Schubert modules $Q(\lambda)$).
Let $\lambda=w\lambda^{+}$ as above. The relative Schubert module $Q(\lambda)$ is set to be the kernel of the (surjective) restriction homomorphism $P(\lambda)\rightarrow{\rm H}^{0}(\partial{X}_{w},\mathcal{L}(\lambda^{+}))$, where the boundary $\partial{X}_{w}$ is the union of the ${X}_{v}$ that are strictly contained in ${X}_{w}$.
2.4 $\mathbf{B}$-cohomological duality
@@
${\mathrm{Ext}}^{i}_{\mathbf{B}}(P(\lambda)^{\ast},Q(\mu))=\left\{\begin{array}[]{l}{\sf k},\quad{\rm for}\quad i=0,\ \lambda=-\mu,\\ 0,\quad{\rm otherwise}.\\ \end{array}\right.$ (4)
See also Theorem 2.9.
@@ 2.5 Excellent order
Recall that we have fixed a Weyl group invariant inner product (,)(~{},~{})( , ) on $X({\mathbf{T}})\otimes_{\mathbb{Z}}\mathbb{R}$.
Definition 2.4.
([48]) Let $\lambda,\mu\in X({\mathbf{T}})$. Define that $\lambda$ is less than $\mu$ in the excellent order, notation $\lambda\leq_{e}\mu$, if either $(\lambda,\lambda)<(\mu,\mu)$ or $\lambda=w\nu$, $\mu=z\nu$ for some $\nu\in X({\mathbf{T}})_{+}$, $w,z\in W$ with $w\leq z$ (in the Bruhat order on $W$).
@@ 2.6 Antipodal excellent order
2.7 Two highest weight category structures on ${\mathrm{Rep}}(\mathbf{B})$
@@
([48, Theorem 1.6, (i)]) The category ${\mathrm{Rep}}(\mathbf{B})$ of rational representations of $\mathbf{B}$ is a highest weight category with respect to the excellent order. The $P(\lambda)$ are the costandard modules for this order.
@@
([48, Theorem 1.6, (ii)]) The category ${\mathrm{Rep}}(\mathbf{B})$ of rational representations of $\mathbf{B}$ is a highest weight category with respect to the antipodal excellent order. The $Q(\lambda)$ are the costandard modules for this order.
@@
The terminology costandard is not used in [48]. In the usual terminology nowadays, cf. [22, §2], $P(\lambda)$ is the costandard module for $<_{e}$ with socle ${\sf k}_{\lambda}$ and $Q(\lambda)$ is the costandard module for $<_{a}$ with socle ${\sf k}_{\lambda}$. Dually, $P(\lambda)^{*}$ is the standard module for $<_{a}$ with head ${\sf k}_{\lambda}^{*}={\sf k}_{-\lambda}$ and $Q(\lambda)^{*}$ is the standard module for $<_{e}$ with head ${\sf k}_{\lambda}^{*}={\sf k}_{-\lambda}$. So equation (4) is about the interaction of standard and costandard modules for $<_{a}$, as in [23, Proposition 3.12].
2.8 Cohomology vanishing for $\mathbf{B}$-modules
@@
([48, Theorem 2.20(i)], [49, Theorem 3.2.6]) Let $\lambda,\mu\in X({\mathbf{T}})$. Then $H^{p}(\mathbf{B},P(\lambda)\otimes Q(\mu))=0$ for $p>0$.
This will be subsumed by Corollary 2.21.
2.9 Filtrations on $\mathbf{B}$-modules
We recall here the fundamental statements concerning filtrations on $\mathbf{B}$-modules. As we emphasized in Section 1.5.1, one has to take some care when citing the literature, because our $\mathbf{B}$ is not $\mathbf{B}^{+}$. We have to convert to our conventions. That means that dominant often becomes antidominant and vice versa.
We begin with recalling the definition of good filtration on a $\mathbf{G}$-module, [20].
Definition 2.10.
A rational $\mathbf{G}$-module $M$ is said to admit a good filtration provided that there exists a finite or infinite increasing filtration
such that each $M_{i}/M_{i-1}=\nabla_{\lambda_{i}}$ for some $\lambda_{i}\in X(\mathbf{T})_{+}$.
Aside 2.11.
An unfortunate side effect of this definition is that if $M$ admits a good filtration, its dimension is at most countable, even if $\mathbf{G}$ acts trivially on $M$. This makes results like Theorem 2.24 a bit complicated. If one wants to keep representation theory clean, it is better to use [49, Definition 4.1.1], which is similar to the next definition.
Definition 2.12.
[49, Definition 2.3.6] A $\mathbf{B}$-module $M$ is said to have an excellent filtration if and only if there exists a filtration $0\subset F_{0}\subset F_{1}\subset\dots$ by $\mathbf{B}$-modules such that $\cup F_{i}=M$ and $F_{i}/F_{i-1}=\oplus P(\lambda_{i})$ for some $\lambda_{i}\in X(\mathbf{T})$. Here $\oplus$ stands for any number of copies, ranging from zero copies to infinitely many.
Remark 2.13.
Definition 2.14.
[49, Definition 2.3.8] A $\mathbf{B}$-module $M$ is said to have a relative Schubert filtration if and only if there exists a filtration $0\subset F_{0}\subset F_{1}\subset\dots$ by $\mathbf{B}$-modules such that $\cup F_{i}=M$ and $F_{i}/F_{i-1}=\oplus Q(\lambda_{i})$ for some $\lambda_{i}\in X(\mathbf{T})$. Here $\oplus$ stands for any number of copies, ranging from zero copies to infinitely many.
Remark 2.15.
Definition 2.16.
A weight $\lambda$ of a $\mathbf{B}$-module $M$ is called extremal if all other weights $\mu$ of $M$ are shorter or equally long: $(\mu,\mu)\leq(\lambda,\lambda)$.
Proposition 2.17.
[49, Proposition 2.3.11] For any $\mathbf{B}$-invariant closed subset $S\subset\mathbf{G}/\mathbf{B}$ and $\lambda\in X(\mathbf{T})_{+}$, the module $H^{0}(S,\mathcal{L}(\lambda))$ has a relative Schubert filtration with layers $Q(w\lambda)$. Moreover, $Q(w\lambda)$ occurs if and only if $w\lambda$ is an extremal weight of $H^{0}(S,\mathcal{L}(\lambda))$ and then has multiplicity one.
@@ 2.9 Filtrations on $\mathbf{B}$-modules
[49, Corollary 2.3.12] The modules ${{\sf H}}_{w}(\lambda)$ have relative Schubert filtration for all $w\in W$ and $\lambda\in X(\mathbf{T})_{+}$.
Theorem 2.19.
([49, Corollary 5.2.3], [39, §5, Corollary 2]) Let $\lambda\in X(\mathbf{T})_{+}$ be a regular dominant weight and $\mu\in X(\mathbf{T})$ be an arbitrary weight. Then $\lambda\otimes Q(\mu)$ has excellent filtration.
Theorem 2.20.
[49, Corollary 5.2.7] Let $\lambda\in X({\mathbf{T}})$ and $\mu\in X({\mathbf{T}})_{+}$. Then $P(\lambda)\otimes\nabla_{\mu}$ has excellent filtration. Note that $\nabla_{\mu}=P(w_{0}\mu)$.
Proof.
This relies on the main results of [39]. For completeness we recall the argument. Let $\lambda^{+}$ be the dominant weight in the Weyl group orbit of $\lambda$ and let $w$ be minimal so that $\lambda=w\lambda^{+}$. By repeated application of the Tensor Identity [32, I Proposition 4.8], we have $P(\lambda)\otimes\nabla_{\mu}={{\sf H}}_{w}({\sf k}_{\lambda^{+}}\otimes\nabla_{\mu})$. Now $k_{\lambda^{+}}\otimes\nabla_{\mu}$ has excellent filtration by [39, §5, Corollary 1]. The result in the Theorem thus follows from [48, Lemma 2.11]. ∎
The next result is a key ingredient in our proofs.
Corollary 2.21.
Let $\lambda,\mu\in X(\mathbf{T})$, $\nu\in X(\mathbf{T})_{+}$. Then $H^{p}(\mathbf{B},P(\lambda)\otimes Q(\mu)\otimes\nabla_{\nu})=0$ for $p>0$.
@@
The vanishing statement of Corollary 2.21 exhibits a remarkable interplay among the three highest weight category structures: one on the category ${\mathrm{Rep}}(\mathbf{G})$ from Section 2.1, and the other two on ${\mathrm{Rep}}(\mathbf{B})$ from Section 2.7. Note also that ${\sf k}=P(0)=Q(0)=\nabla(0)$ is costandard in all three highest weight category structures.
2.10 Cohomological criteria
Theorem 2.23.
[20, Theorem 1.1’] If $\lambda\in X(\mathbf{T})_{+}$ and $M$ is a $\mathbf{G}$-module with a good filtration then ${\rm H}^{i}(\mathbf{G},\nabla_{\lambda}\otimes M)=0$ for $i>0$.
Theorem 2.24.
[27, Theorem 4, Corollary 7] Let $M$ be a finitely/countably generated rational $\mathbf{G}$-module satisfying ${\rm H}^{1}(\mathbf{G},\nabla_{\mu}\otimes M)=0$ for all $\mu\in X(\mathbf{T})_{+}$. Then $M$ admits a good filtration.
Theorem 2.25.
[32, II Proposition 4.21, B.9] Let $\lambda$, $\mu\in X(\mathbf{T})_{+}$. Then $\nabla_{\lambda}\otimes\nabla_{\mu}$ admits a good filtration.
Theorem 2.26.
[48, Theorem 3.2.7] Let $M$ be a $\mathbf{B}$-module such that for every $\lambda\in X(\mathbf{T})$, $H^{1}(\mathbf{B},M\otimes Q(\lambda))=0$. Then, $M$ has an excellent filtration.
2.11 Forms of the $P(\lambda)$, $Q(\mu)$ over the integers
For the results over ${\mathbb{Z}}$ we will use ${\mathbb{Z}}$-forms $P(\lambda)_{\mathbb{Z}}$, $Q(\mu)_{\mathbb{Z}}$ of $P(\lambda)$, $Q(\mu)$ respectively. They are constructed and studied in [49, Chapter 7]. Both $P(\lambda)_{\mathbb{Z}}$ and $Q(\mu)_{\mathbb{Z}}$ are finitely generated and free over ${\mathbb{Z}}$. They share many properties with their counterparts over fields. We often drop the ${\mathbb{Z}}$ from the notation $P(\lambda)_{\mathbb{Z}}$, $Q(\mu)_{\mathbb{Z}}$. One could follow [23, section 3] to discuss in what sense one has highest weight category structures on the exact category of $\mathbb{B}$-modules that are finitely generated and free over ${\mathbb{Z}}$. Instead we will invoke the Universal coefficient Theorem for cohomology to move back and forth between the case that ${\sf k}$ is a field and the case ${\sf k}={\mathbb{Z}}$. We need that tool anyway.
3 Rappels: Weyl groups
3.1 The Demazure product
Let $(W,S)$ be the Coxeter presentation of the Weyl group of a simple simply connected group $\mathbf{G}$ of rank $r$. The Demazure product on $W$ is defined by replacing the relation $s_{\alpha}\cdot s_{\alpha}=e$ for a simple root $\alpha$ with $s_{\alpha}\star s_{\alpha}=s_{\alpha}$. Equivalently, for any $w\in W$ and any simple root $\alpha$, define $w\star s_{\alpha}$ to be the longer of $ws_{\alpha}$ or $w$. A sequence of simple reflections is reduced if and only if the usual product $s_{i}s_{2}\dots s_{k}$ is equal to the Demazure product $s_{i}\star\dots\star s_{{k}}$. If $v,w\in W$, then $\overline{\mathbf{B}v\mathbf{B}}~{}\overline{\mathbf{B}w\mathbf{B}}=\overline{\mathbf{B}(v\star w)\mathbf{B}}$.
Lemma 3.1.
If $v,w\in W$, then ${{\sf H}}_{v}\circ{{\sf H}}_{w}={{\sf H}}_{v\star w}$.
Proof.
See [14, 1.3., Corollary and Theorem 3.1] or [49, Proposition 2.2.5]. ∎
3.2 Total order $\prec$ on $W$, refining the Bruhat order
@@
The order $\prec$ gives an order on the Steinberg basis, often in conflict with $<_{a}$ and $>_{e}$.
3.3 Combinatorics
First we recall some facts from [30] about $W$ and its action on $X(\mathbf{T})$. As usual $w_{0}$ denotes the longest element of $W$.
Lemma 3.3.
Let $v,w\in W$. The following are equivalent
  • $v<w$
  • $v^{-1}<w^{-1}$
  • $w_{0}w<w_{0}v$
  • $ww_{0}<vw_{0}$.
Proof.
For the first three, see [30, page 119]. Use that $(ww_{0})^{-1}=w_{0}w^{-1}$, $(vw_{0})^{-1}=w_{0}v^{-1}$. ∎
Lemma 3.4.
Proof.
See [30, page 116]. ∎
Lemma 3.5.
Proof.
See [30, page 16]. We usually prefer to multiply from the left. ∎
Lemma 3.6.
Let $w,z\in W$ with $\ell(wz)=\ell(w)+\ell(z)$. Then $w\leq w_{0}z^{-1}$.
Proof.
Lemma 3.7.
Every $w\in W$ may be written uniquely as $vz$, where $z\in W_{I}$ and $v\in W^{I}$ is the minimal coset representative of the coset $wW_{I}$. The assignment of $v$ to $w$ respects the Bruhat order.
Proof.
Compare [30, 1.10, 1.15]. For comparison with [3, Section 3], note that, for $v\in W^{I}$, $\ell(v)$ is the number of reflecting hyperplanes that separate $v\lambda$ from $\lambda$. ∎
Recall that we denote the Demazure product on $W$ by $\star$.
Lemma 3.8.
Lemma 3.9.
Let $x$, $y$, $s\in W$ with $s$ simple and $x\leq y\star s$. Choose $v$ minimal so that $v\star s=x\star s$. Then $v\leq y$.
Proof.
Note that $v$ is the minimal representative of the coset $x\langle s\rangle$. If $w$ is the minimal representative of the coset $y\langle s\rangle$, then $v\leq w\leq y$. ∎
Lemma 3.10.
Proof.
4 $\mathbf{T}$-equivariant ${\sf K}$-theory of $\mathbf{G}/\mathbf{B}$